Each ounce of substance A supplies is 4% of nutrition patient needs. Substance B is 10% and C is 14%. Restrictions require A & C to be given in equal amounts, the amount of B is 1/5 of either A or C, find the number of ounces each substance should be to provide 100% nutrition.
Answer by ankor@dixie-net.com(22740)

Each ounce of substance A supplies is 4% of nutrition patient needs.
Substance B is 10% and C is 14%.
Restrictions require A & C to be given in equal amounts, the amount of B is 1/5 of either A or C, find the number of ounces each substance should be to provide 100% nutrition.
:
let a = amt of substance A and amt of substance C
let .2a = amt of substance B
:
4A + 10B + 14C = 100
4a + 10(.2a) + 14a = 100
4a + 2a + 14a = 100
20a = 100
a = 100/20
a = 5 oz of Substance A and 5 oz of substance C
.2(5) = 1 oz of substance B